# Probabilistic Guarantees to Explicit Constructions: Local Properties of Linear Codes

### v1

@@ 5 Constructions for Local Properties
    monotone-increasing, and
  3. (iii)
    independent of coordinate permutations.
A local code property, informally speaking, is defined by the inclusion of bad sets of vectors. A monotone-increasing code property is one for which the following is true: if $\mathcal{C}$ is in $\mathcal{P}$, then every $\mathcal{C}^{\prime}$ for which $\mathcal{C}^{\prime}\supseteq\mathcal{C}$ holds, also lies in $\mathcal{P}$. We say that a property $\mathcal{P}$ is independent of coordinate permutations if for every code $\mathcal{C}\in\mathcal{P}$, $\mathcal{C}^{\prime}$ also lies in $\mathcal{P}$, where $\mathcal{C}^{\prime}$ is obtained by applying a permutation on the coordinates to every codeword in $\mathcal{C}$. An example of local, monotone-increasing code property that is independent of coordinate permutations is the complement of $(\rho,L)$-list-decodability.
@@
Our first modification is to define local profiles as tuples of matrices, instead of subspaces. We then require that in order for a matrix $A$ to be contained in a local profile, each row of $A$ should lie in the kernel of the matrix corresponding to the row index. This is necessary in order to address the problem of differing field sizes. Recall that the alphabet of $\mathcal{C}_{\textrm{AEL}}$ is $\Sigma_{\textrm{in}}^{d}$, where $\Sigma_{\textrm{in}}$ is the alphabet of the inner code. Thus, if our inner code is over a field ${\mathbb{F}}_{q}$, one can view the alphabet of $\mathcal{C}_{\textrm{AEL}}$ as being equal to ${\mathbb{F}}_{q}^{d}$. Note that one can naturally view the vectors in ${\mathbb{F}}_{q}^{d}$ as elements in the extension field ${\mathbb{F}}_{Q}$, where $Q=q^{d}$. This fact ensures that the rows of matrices in ${\mathbb{F}}_{q}^{L\times L}$, when viewed as linear constraints, will be applicable to vectors in ${\mathbb{F}}_{q}^{L}$ and ${\mathbb{F}}_{Q}^{L}$ simultaneously, as ${\mathbb{F}}_{Q}$ is an extension field of ${\mathbb{F}}_{q}$.
@@ 5.1 Preliminaries
Fix a locality parameter $L\in{\mathbb{N}}$, and a finite field ${\mathbb{F}}_{q}$. Throughout this section, we restrict attention to $\mathbb{F}_{q}$-linear codes. Accordingly, the term “linear” will henceforth always refer to $\mathbb{F}_{q}$-linear.
@@
where for each $t\in[T]$, we have $f_{t}\in[0,1]$ and $\sum_{t\in[T]}f_{t}=1$. The matrices $\mathbf{M}_{t}$ are not required to be pairwise distinct.
@@
where $M_{i}\in{\mathbb{F}}_{q}^{L\times L}$ for each $i\in[n]$. Moreover, as prescribed by $\mathcal{V}$, for each $t\in[T]$, the matrix $\mathbf{M}_{t}$ appears in exactly $f_{t}\cdot n$ coordinates in $M_{n}(\mathcal{V})$.
We emphasize that $M_{n}(\mathcal{V})$ is not unique given a local profile description $\mathcal{V}$. In fact, it is easily seen that all permutations of the entries of $M_{n}(\mathcal{V})$ are valid local profiles that can be created using $\mathcal{V}$. We will denote the set of all local profiles that can be created from $\mathcal{V}$ (for a block length $n$) by $\mathcal{V}_{n}$.
@@
At this point we provide a brief discussion about the LCL properties $\mathcal{P}$ considered henceforth. In general, we allow the fractions $f_{1},\ldots,f_{T}$ to range over sub-intervals of $[0,1]$, and regard all local profile descriptions with such fractions as belonging to $\mathcal{P}$. Consequently, $\mathcal{P}$ is uncountably infinite. This, however, does not pose a difficulty, because for a block length $n$, the set of local profile descriptions satisfying $\left|\mathcal{V}_{n}\right|>0$ is itself finite, as by definition, each fraction in such local profile descriptions must be an exact multiple of $1/n$. Furthermore, for any such local profile description, the corresponding set of local profiles $\mathcal{V}_{n}$ is also finite. Specifically, $\left|\mathcal{V}_{n}\right|$ is equal to the number of ways we can arrange $n$ objects in a row, where we have $f_{t}\cdot n$ objects of type $t$, for $t\in T$. It follows, therefore, that
@@
Definition 5.18.
We say an $L$-LCL property $\mathcal{P}$ is reasonable if for every block length $n$, the total number of associated local profiles is at most
$\sum_{\mathcal{V}\in\mathcal{P}}\left|\mathcal{V}_{n}\right|\leq q^{\kappa_{q}(\mathcal{P})\cdot n},$
where the term $\kappa_{q}(\mathcal{P})$ approaches 0 as $q\rightarrow\infty$.
We note that $\kappa_{q}(\mathcal{P})$ is allowed to depend on $L$ in an arbitrary fashion, although a worse dependence would result in $q$ being large (with respect to $L$) as well. Most local properties considered in the literature, such as list-decoding, list-recovery and perfect hash matrices are reasonable local properties.
@@
$\operatorname*{\arg\!\max}\left\{\mathsf{\Phi}(\mathcal{V},*,R)\right\}:=\left\{W\in\mathcal{L}({\mathbb{F}}_{q}^{L})\mid\mathsf{\Phi}(\mathcal{V},U,R)=\max_{U\in\mathcal{L}({\mathbb{F}}_{q}^{L})}\mathsf{\Phi}(\mathcal{V},U,R))\right\}.$
@@ 5.2 Implied Local Profile Descriptions
We now give some intuition for Definition˜5.21.
Observation 5.22.
@@
Let $W_{\mathcal{V}}$ be the canonical subspace guaranteed by ˜5.20 for $\mathcal{V}$. Recall that for all $R\leq R_{\mathcal{V}}$ and for every $U\in\mathcal{L}({\mathbb{F}}_{q}^{L})$ satisfying $\dim U\geq\dim W_{\mathcal{V}}$, $W_{\mathcal{V}}$ satisfies
@@
By the guarantee for $W_{\mathcal{V}}$ (see ˜5.20), we have for all subspaces $U\in\mathcal{L}({\mathbb{F}}_{q}^{L})$ satisfying $\dim U\geq\dim W_{\mathcal{V}}$,
@@
where Eq.˜11 follows from ˜5.27, stated and proved below. Moreover, Eq.˜12 follows from an application of the rank-nullity theorem. The final inequality follows from Eq.˜10, owing to the fact that $U\supset W_{\mathcal{V}}$, given that $U^{\prime}$ is a non-trivial subspace. ∎
@@
Let $W_{\mathcal{V}},\psi_{\mathcal{V}},U^{\prime}$ and $U$ be as defined in Lemma˜5.26. Then, for every subspace K in ${\mathbb{F}}_{q}^{L}$, we have
@@
By Lemma˜5.28, we see that for an $L$-LCL property $\mathcal{P}$, the total number of local profiles that can be created according to local profile descriptions from the set $\bigcup_{\mathcal{V}\in\mathcal{P}}\operatorname{Rob}_{\Delta}(\mathcal{V}^{\mathsf{imp}})$ is at most
@@
where the inequality follows from Lemma˜5.26. The result follows by union bounding over all non-zero subspaces $U$, and all local profiles created according to local profile descriptions from the set $\bigcup_{\mathcal{V}\in\mathcal{P}}\operatorname{Rob}_{\Delta}(\mathcal{V}^{\mathsf{imp}})$. ∎
We now turn to analyze the upper bound on the probability in Lemma˜5.29. The bound is non-trivial if and only if $\sum_{\mathcal{V}\in\mathcal{P}}|\mathcal{V}_{n}|\cdot 2^{n}\cdot q^{L^{2}-\varepsilon n}\cdot q^{\Delta\cdot Ln}<1$. For reasonable $L$-LCL properties, this is achieved when
$\kappa_{q}(\mathcal{P})+\log_{q}2+\frac{L^{2}}{n}-\varepsilon+\Delta\cdot L<0.$
@@
Here, we are interested in the smallest prime power $q$ that satisfies the above inequality. Such a $q$ exists, since $\lim_{q\rightarrow\infty}(\kappa_{q}(\mathcal{P})+\log_{q}2)=0$.
@@
For a reasonable $L$-LCL property $\mathcal{P}$ and $R\leq R_{\mathcal{P}}-\varepsilon$, there exists a minimum prime power $q$ that is solely a function of $\varepsilon,L,$ and $\kappa_{q}(\mathcal{P})$, such that the following holds: There exists a linear code $\mathcal{C}\subseteq{\mathbb{F}}_{q}^{n}$ for every block length $n\geq 4L^{2}/\varepsilon$ that contains no non-zero matrices satisfying local profile descriptions from the set $\bigcup_{\mathcal{V}\in\mathcal{P}}\operatorname{Rob}_{\Delta}(\mathcal{V}^{\mathsf{imp}})$.
@@
For a block length $n\geq 4L^{2}/\varepsilon$ and a reasonable $L$-LCL property $\mathcal{P}$, let $\mathcal{C}$ be a linear code $\mathcal{C}$ as described in ˜5.30. Then,
@@
The number of non-zero matrices satisfying local profile descriptions from the set $\bigcup_{\mathcal{V}\in\mathcal{P}}\operatorname{Rob}_{\Delta}(\mathcal{V}^{\mathsf{imp}})$ is at most the total number of matrices of dimensions $n\times L$, which is equal to $q^{nL}$. Given the parity check matrix of a linear code $\mathcal{C}\subseteq{\mathbb{F}}_{q}^{n}$, we can check whether every column of a matrix $A\in{\mathbb{F}}_{q}^{n\times L}$ lies in $\mathcal{C}$ in time at most $n^{2}L$. Therefore, for each linear code, we can perform the aforementioned check in time at most $n^{2}L\cdot q^{nL}$. Consequently, the total runtime is at most
@@
We now prove that $\mathcal{C}$ does not satisfy $\mathcal{P}$. Assume for contradiction that it does. Then, there is a $\mathcal{V}\in\mathcal{P}$ such that $\mathcal{C}$ contains $\mathcal{V}$. By ˜5.22, $\mathcal{C}$ contains a matrix $B$ that satisfies $\mathcal{V}^{\mathsf{imp}}$. The matrix $B$ was obtained from some matrix $A\subseteq\mathcal{C}$ having pairwise distinct columns, by applying a map $\psi$ to every row of $A$. By the definition of $\mathcal{V}^{\mathsf{imp}}$, there exist $i,j\in[L]$, where $i\neq j$, such that the kernel of $\psi$ consists of a subspace whose vectors agree at coordinates $i,j$. This implies that $B$ is non-zero. Furthermore, by ˜5.25, $B$ also satisfies every $\mathcal{V}^{\mathsf{imp}}_{\Delta}\in\operatorname{Rob}_{\Delta}(\mathcal{V})$. This contradicts the fact that $\mathcal{C}$ is of the form as specified in ˜5.30. ∎
Recall that the code produced by the AEL construction, $\mathcal{C}_{\textrm{AEL}}$, has alphabet $\Sigma_{\textrm{in}}^{d}$, where $\Sigma_{\textrm{in}}$ is the alphabet of the inner code. We take the inner code to be over ${\mathbb{F}}_{q}$, and therefore interpret $\mathcal{C}_{\textrm{AEL}}$ as a code over the extension field ${\mathbb{F}}_{Q}$, where $Q=q^{d}$. Furthermore, by inspection, the definitions of containment and satisfiability (Definition˜5.3, Definition˜5.4, Definition˜5.5, Definition˜5.7) also apply to matrices and codes over ${\mathbb{F}}_{Q}$. In particular, they apply to $\mathcal{C}_{\textrm{AEL}}\subseteq{\mathbb{F}}_{Q}^{N}$.

### v2

@@ 5 Constructions for Local Properties
    monotone-increasing.
A local code property, informally speaking, is defined by the inclusion of bad sets of vectors. A monotone-increasing code property is one for which the following is true: if $\mathcal{C}$ is in $\mathcal{P}$, then every $\mathcal{C}^{\prime}$ for which $\mathcal{C}^{\prime}\supseteq\mathcal{C}$ holds, also lies in $\mathcal{P}$. An example of local, monotone-increasing code property is the complement of $(\rho,L)$-list-decodability.
@@
Our first modification is to define local profiles as tuples of matrices, instead of subspaces. We then require that in order for a matrix $A$ to be contained in a local profile, each row of $A$ should lie in the kernel of the matrix corresponding to the row index. This is necessary in order to address the problem of differing field sizes. Recall that the alphabet of $\mathcal{C}_{\textnormal{AEL}}$ is $\Sigma_{\textnormal{in}}^{d}$, where $\Sigma_{\textnormal{in}}$ is the alphabet of the inner code. Thus, if our inner code is over a field ${\mathbb{F}}_{q}$, one can view the alphabet of $\mathcal{C}_{\textnormal{AEL}}$ as being equal to ${\mathbb{F}}_{q}^{d}$. Note that one can naturally view the vectors in ${\mathbb{F}}_{q}^{d}$ as elements in the extension field ${\mathbb{F}}_{Q}$, where $Q=q^{d}$. This fact ensures that the rows of matrices in ${\mathbb{F}}_{q}^{L\times L}$, when viewed as linear constraints, will be applicable to vectors in ${\mathbb{F}}_{q}^{L}$ and ${\mathbb{F}}_{Q}^{L}$ simultaneously, as ${\mathbb{F}}_{Q}$ is an extension field of ${\mathbb{F}}_{q}$.
@@ 5.1 Preliminaries
Fix a locality parameter $L\in{\mathbb{N}}$. Additionally, consider finite fields $\mathbb{F}_{q_{0}},\mathbb{F}_{q}$ where $q_{0},q$ are powers of the same prime that satisfy $q_{0}\leq q$. It follows that $\mathbb{F}_{q}$ is an extension field of $\mathbb{F}_{q_{0}}$. Throughout the paper, we fix a $q_{0}$ and define LCL properties using elements from $\mathbb{F}_{q_{0}}$, and work with ${\mathbb{F}}_{q}$-linear codes over the field $\mathbb{F}_{q}$. The precise value of $q$ will be fixed later.
Accordingly, the term “linear” will henceforth always refer to $\mathbb{F}_{q}$-linear.
@@
where for each $t\in[T]$, we have $f_{t}\in[0,1]$, $\sum_{t\in[T]}f_{t}=1$, and $\mathbf{M}_{t}\in\mathbb{F}_{q_{0}}^{L\times L}$. The matrices $\mathbf{M}_{t}$ are not required to be pairwise distinct.
@@
where $M_{i}\in\mathbb{F}_{q_{0}}^{L\times L}$ for each $i\in[n]$. Moreover, as prescribed by $\mathcal{V}$, for each $t\in[T]$, the matrix $\mathbf{M}_{t}$ appears in exactly $f_{t}\cdot n$ coordinates in $M_{n}(\mathcal{V})$.
We emphasize that $M_{n}(\mathcal{V})$ is not unique for a local profile description $\mathcal{V}$. In fact, it is easily seen that all permutations of the entries of $M_{n}(\mathcal{V})$ are valid local profiles that can be created using $\mathcal{V}$. We will denote the set of all local profiles that can be created from $\mathcal{V}$ (for a block length $n$) by $\mathcal{V}_{n}$.
@@
At this point we provide a brief discussion about the LCL properties $\mathcal{P}$ considered henceforth. In general, we allow the fractions $f_{1},\ldots,f_{T}$ to range over some fixed sub-intervals of $[0,1]$, and regard all local profile descriptions with such fractions as belonging to $\mathcal{P}$. Consequently, $\mathcal{P}$ is uncountably infinite. This, however, does not pose a difficulty, because for a block length $n$, the set of local profile descriptions satisfying $\left|\mathcal{V}_{n}\right|>0$ is itself finite, as by definition, each fraction in such local profile descriptions must be an exact multiple of $1/n$. Furthermore, for any such local profile description, the corresponding set of local profiles $\mathcal{V}_{n}$ is also finite. Specifically, $\left|\mathcal{V}_{n}\right|$ is equal to the number of ways we can arrange $n$ objects in a row, where we have $f_{t}\cdot n$ objects of type $t$, for $t\in T$. It follows, therefore, that
@@
Since our definition of LCL properties is derived from the notion of local profile descriptions, the properties we consider are coordinate-permutation invariant. An LCL property $\mathcal{P}$ is said to be coordinate-permutation invariant if, for any local profile $M=(M_{1},\ldots,M_{n})$ associated with $\mathcal{P}$, all local profiles obtained by permuting the order of the matrices in $M$ are also associated with $\mathcal{P}$.
At first glance, this might appear to restrict the class of properties we can consider; however, it does not. Indeed, for a local profile description $\mathcal{V}$, the threshold rate $R_{\mathcal{V}}$ depends only on the frequency of the matrices appearing in the description, and not on their order. Consequently, the threshold rate remains the same whether we consider a single local profile described by $\mathcal{V}$ or all of them. Thus, if we initially intended to include in our LCL property $\mathcal{P}$ only a subset of the local profiles described by $\mathcal{V}$, we may in fact include all of them without any loss in threshold rate. This does result in an increase in alphabet size, though it remains constant.
@@
$\operatorname*{\arg\!\max}\left\{\mathsf{\Phi}(\mathcal{V},*,R)\right\}:=\left\{W\in\mathcal{L}({\mathbb{F}}_{q}^{L})\mid\mathsf{\Phi}(\mathcal{V},W,R)=\max_{U\in\mathcal{L}({\mathbb{F}}_{q}^{L})}\mathsf{\Phi}(\mathcal{V},U,R)\right\}.$
@@ 5.2 Implied Local Profile Descriptions
We now give some intuition for Definition˜5.20.
Observation 5.21.
@@
Let $W_{\mathcal{V}}$ be the canonical subspace guaranteed by ˜5.19 for $\mathcal{V}$. Recall that for all $R\leq R_{\mathcal{V}}$ and for every $U\in\mathcal{L}({\mathbb{F}}_{q}^{L})$ satisfying $\dim U\geq\dim W_{\mathcal{V}}$, $W_{\mathcal{V}}$ satisfies
@@
By the guarantee for $W_{\mathcal{V}}$ (see ˜5.19), we have for all subspaces $U\in\mathcal{L}({\mathbb{F}}_{q}^{L})$ satisfying $\dim U\geq\dim W_{\mathcal{V}}$,
@@
where Eq.˜11 follows from ˜5.26, stated and proved below. Moreover, Eq.˜12 follows from an application of the rank-nullity theorem. The final inequality follows from Eq.˜10, owing to the fact that $U\supset W_{\mathcal{V}}$, given that $U^{\prime}$ is a non-trivial subspace. ∎
@@
Let $W_{\mathcal{V}},\psi_{\mathcal{V}},U^{\prime}$ and $U$ be as defined in Lemma˜5.25. Then, for every subspace K in ${\mathbb{F}}_{q}^{L}$, we have
@@
By Lemma˜5.27, we see that for an $L$-LCL property $\mathcal{P}$, the total number of local profiles that can be created according to local profile descriptions from the set $\bigcup_{\mathcal{V}\in\mathcal{P}}\operatorname{Rob}_{\Delta}(\mathcal{V}^{\mathsf{imp}})$ is at most
@@
where the inequality follows from Lemma˜5.25. The result follows by union bounding over all non-zero subspaces $U$, and all local profiles created according to local profile descriptions from the set $\bigcup_{\mathcal{V}\in\mathcal{P}}\operatorname{Rob}_{\Delta}(\mathcal{V}^{\mathsf{imp}})$. ∎
We now turn to analyze the upper bound on the probability in Lemma˜5.28.
Definition 5.29.
For a prime power $q$ and all sufficiently large $n$, we define $\kappa_{q}(\mathcal{P})\in\mathbb{R}$ as
$\kappa_{q}(\mathcal{P}):=\inf_{\kappa}\left\{\kappa:\sum_{\mathcal{V}\in\mathcal{P}}|\mathcal{V}_{n}|\leq q^{\kappa n}\right\}.$
The probability upper bound is non-trivial if and only if $\sum_{\mathcal{V}\in\mathcal{P}}|\mathcal{V}_{n}|\cdot 2^{n}\cdot q^{L^{2}-\varepsilon n}\cdot q^{\Delta\cdot Ln}<1$. This is achieved when
$\kappa_{q}(\mathcal{P})+\log_{q}2+\frac{L^{2}}{n}-\varepsilon+\Delta\cdot L<0,$
@@
Here, we are interested in the smallest prime power $q$ that satisfies the above inequality. Such a $q$ exists. Indeed, this is seen because the total number of local profiles associated with $\mathcal{P}$—equal to $q^{\kappa_{q}(\mathcal{P})n}$—is at most $q_{0}^{L^{2}n}$. Recall that $\mathbb{F}_{q_{0}}$ is the field over which constraints of $\mathcal{P}$ were defined. Thus, $\kappa_{q}(\mathcal{P})\leq\log_{q}(q_{0})\cdot L^{2}$, which implies
$\lim_{q\rightarrow\infty}(\kappa_{q}(\mathcal{P})+\log_{q}2)=0.$
We choose the smallest $q$ so that $\kappa_{q}(\mathcal{P})+\log_{q}2$ is lesser than $\varepsilon/4$, and thus Eq.˜13 is satisfied.
@@
For an $L$-LCL property $\mathcal{P}$ and $R\leq R_{\mathcal{P}}-\varepsilon$, there exists a minimum prime power $q$ that is solely a function of $\varepsilon,L,q_{0},$ and $\mathcal{P}$, such that the following holds: There exists a linear code $\mathcal{C}\subseteq{\mathbb{F}}_{q}^{n}$ for every block length $n\geq 4L^{2}/\varepsilon$ that contains no non-zero matrices satisfying local profile descriptions from the set $\bigcup_{\mathcal{V}\in\mathcal{P}}\operatorname{Rob}_{\Delta}(\mathcal{V}^{\mathsf{imp}})$.
@@
For a block length $n\geq 4L^{2}/\varepsilon$ and an $L$-LCL property $\mathcal{P}$, let $\mathcal{C}$ be a linear code $\mathcal{C}$ as described in ˜5.30. Then,
@@
The number of non-zero matrices satisfying local profile descriptions from the set $\bigcup_{\mathcal{V}\in\mathcal{P}}\operatorname{Rob}_{\Delta}(\mathcal{V}^{\mathsf{imp}})$ is at most the total number of matrices of dimensions $n\times L$, which is equal to $q_{0}^{nL}\leq q^{nL}$. Given the parity check matrix of a linear code $\mathcal{C}\subseteq{\mathbb{F}}_{q}^{n}$, we can check whether every column of a matrix $A\in{\mathbb{F}}_{q}^{n\times L}$ lies in $\mathcal{C}$ in time at most $n^{2}L$. Therefore, for each linear code, we can perform the aforementioned check in time at most $n^{2}L\cdot q^{nL}$. Consequently, the total runtime is at most
@@
We now prove that $\mathcal{C}$ does not satisfy $\mathcal{P}$. Assume for contradiction that it does. Then, there is a $\mathcal{V}\in\mathcal{P}$ such that $\mathcal{C}$ contains $\mathcal{V}$. By ˜5.21, $\mathcal{C}$ contains a matrix $B$ that satisfies $\mathcal{V}^{\mathsf{imp}}$. The matrix $B$ was obtained from some matrix $A\subseteq\mathcal{C}$ having pairwise distinct columns, by applying a map $\psi$ to every row of $A$. By the definition of $\mathcal{V}^{\mathsf{imp}}$, there exist $i,j\in[L]$, where $i\neq j$, such that the kernel of $\psi$ consists of a subspace whose vectors agree at coordinates $i,j$. This implies that $B$ is non-zero. Furthermore, by ˜5.24, $B$ also satisfies every $\mathcal{V}^{\mathsf{imp}}_{\Delta}\in\operatorname{Rob}_{\Delta}(\mathcal{V})$. This contradicts the fact that $\mathcal{C}$ is of the form as specified in ˜5.30. ∎
Recall that the code produced by the AEL construction, $\mathcal{C}_{\textnormal{AEL}}$, has alphabet $\Sigma_{\textnormal{in}}^{d}$, where $\Sigma_{\textnormal{in}}$ is the alphabet of the inner code. We take the inner code to be over ${\mathbb{F}}_{q}$, and therefore interpret $\mathcal{C}_{\textnormal{AEL}}$ as a code over the extension field ${\mathbb{F}}_{Q}$, where $Q=q^{d}$. Furthermore, by inspection, the definitions of containment and satisfiability (Definition˜5.3, Definition˜5.4, Definition˜5.5, Definition˜5.7) also apply to matrices and codes over ${\mathbb{F}}_{Q}$. In particular, they apply to $\mathcal{C}_{\textnormal{AEL}}\subseteq{\mathbb{F}}_{Q}^{N}$.